OpenStudy (nuccioreggie):

Which statement is true about the equation fraction 3 over 4z = fraction 1 over 4z − 3 + 5? It has no solution. It has one solution It has two solutions. It has infinitely many solutions.

OpenStudy (nuccioreggie):

@whovianchick

OpenStudy (whovianchick):

\[3/4z=1/4z -3+5\]

OpenStudy (whovianchick):

is that the equation?

OpenStudy (nuccioreggie):

yes

OpenStudy (whovianchick):

if so 3/4z = 1/4z -3 +5 3/4z= 1/4z +2 (combine like terms) 2/4z= 2 (subtract 1/4 from both sides) 1/2z=2 (simplify) z=4 (multiply both sides by 2) One solution

OpenStudy (nuccioreggie):

o ok